You are considering an investment in a 40-year security. The security will pay RM25 a year
at the end of each of the first three years. The security will then pay RM30 a year at the end
of each of the next 20 years. The nominal interest rate is assumed to be 8 percent, and the
current price (present value) of the security is RM360.39. Given this information, what is
the equal annual payment to be received from Year 24 through Year 40 (i.e., for 17 years)?
1
Expert's answer
2016-05-10T08:35:03-0400
T = 40, payments of RM25 a year first three years, RM30 next 20 years, i = 8%, P = RM360.39.
If i = 8% and P = RM360.39, then the total amount of interest for 40 years will be 360.39*40*0.08 = RM1153.25.
First 23 years total amount of payments received will be 25*3 + 30*20 = 675, so the equal annual payment to be received from Year 24 through Year 40 (i.e., for 17 years) is (1153.25 - 675)/17 = RM28.13.